# Towards Autonomous Reinforcement Learning for Real-World Robotic Manipulation with Large Language Models

### v2

@@ IV-D Autonomous Skill Learning: Policy Learning
Figure 3: Top: Success rates for SAC agent training on the three selected tasks. The proposed reward formalization is compared with the performance of the rewards completely defined by GPT-4. For each task, 10 rewards are generated and each reward is tested with 10 different random seeds. Plots show median and 25-75 percentiles. Bottom: The agent performing the three selected tasks. The agent (ball) is purple when holding the object (rectangle), while green otherwise.
V Experiments
In this section, we present the experimental evaluations performed to validate the proposed method. First, we validate the effectiveness of the reward formalization described in Section IV-C on three tasks on a simple 2D simulation. Second, we present the results of the Autonomous RL pipeline for 10 different robotics manipulation tasks, including demonstrations on the real setup.
V-A 2D Environments
To assess the effectiveness of the proposed reward formalization, we implemented a 2D simulation that mimics the simplified dynamics of a robotics manipulation scene. In this environment, the agent controls the movements of a point on a 2D vertical plane, a rectangular object is present on the “floor” of the plane. When in proximity to the object, the agent can perform grasping, which constrains the object to move solidly with the agent, until it stops the grasping, releasing the object. We consider three simple manipulation tasks: (i) a grasp and lift task, (ii) a grasp and slide task, (iii) a placing task. Figure 3 (bottom) shows an example execution for each task in the 2d environment. In the first two tasks, the agent is initialized to a constant position $x_{0}$, while the object is initialized in a position $x_{r}\sim p(x_{r})$. In the third task, the object and the agent are initialized next to each other, randomly at the top of the environment, $(x_{0},x_{r})\sim p(x_{0},x_{r})$. Then the agent must grasp the object and bring it to a target location $P$.
For each task, we run ARCHIE, and the same reward generation without formalization defined in Section IV-C, i.e., GPT-4 can define the reward function without formalization constraints, equivalent to the first step of reward generation of [15, 18] and similar works. For each task, we generate 10 rewards for both reward generations, using the same prompt instructions, except for the part that defines the formalization of the reward. The prompt structure is available is the appendix. We train 10 SAC agents for each reward using different random seeds and monitor the success rates of the trained policies. We use a single training environment ($N=1$) per agent. The results are reported in Figure 3 (top), where we show the statistics of the success rate curves of the agents trained by ARCHIE and the statistics of the rewards generated by GPT-4.
In the training horizons considered, the agents trained by ARCHIE with the proposed formalization, consistently solve all tasks, while the agents trained with the rewards completely formalized by GPT-4 are much less consistent. This suggests that the proposed formalization effectively stabilizes policy learning toward task completion.
V-B Robotic Manipulation
We select 10 robotic manipulation tasks, including pushing, picking, insertion, and dual-arm operations. For each task, we define the natural language task description, as described in Section IV-B, which we use to generate a reward with ARCHIE’s proposed reward generation. The same descriptions are exploited to generate code policies [41] with GPT-4, using the same interface as the RL policies.
Each reward function, including the baseline, is used to train 3 SAC agents, for each task, using different random seeds. Each agent explores $N=4$ environments in parallel, as described in Section IV-D and Algorithm 1.
The training results are reported in Figure 4, where we compare the agents in terms of success rate. The figure includes the average success rates of 10 different code policies for each task. The code policies do not solve any of the selected tasks, while the rewards generated by ARCHIE guide the agents toward task completion in all tasks, reaching high success rates in most tasks. Instead, the distance-based baseline rewards only solve a few of the tasks, with low success rates. These results suggest that the reasoning capabilities of the LLM are adequate to select the reward terms related to the objective task, while practically would struggle to define the low-level skills required to solve these tasks without ad-hoc interfaces. Moreover, a proper mathematical structure in the reward formalization is fundamental to exploit the code generation capabilities of the LLM.
Figure 5 depicts the YuMi robot performing some of the tasks in the lab settings, while videos of the experiments are included in the demonstration video. The robot is controlled with a joint position controller, tracking the trajectories performed by the policy trained in simulation. The gripper is operated with a combination of force/position control.
Figure 5: The YuMi robot performing some of the selected tasks.
VI Discussion and Conclusions
In this work, we proposed ARCHIE, an unsupervised pipeline leveraging GPT-4 to automate reward function generation for Reinforcement Learning (RL) in robotic manipulation. By utilizing Large Language Models (LLMs), our method translates human-readable task descriptions into structured reward functions, enabling training of RL agents in simulated environments with minimal manual intervention. Unlike previous methods, our approach integrates both reward generation and task success criteria formulation, creating a one-shot process for autonomous skill acquisition.
Furthermore, our proposed reward formalization ensures stable policy learning, at the first training attempt, by balancing shaping and terminal rewards, mitigating issues related to reward misalignment. In contrast, other works rely on reward reflection to iteratively improve the generated rewards, thus requiring several RL policy training attempts and prompts to the LLM to converge to a stable policy. In the 2D environments, we showed that the proposed one-shot reward generation is much more effective than the first step of reward generation completely formalized by GPT-4, as done in previous works, such as [15, 18].
While our approach significantly advances automated RL training, several avenues remain unexplored for future research. The introduction of reward function refinement through real-world feedback and human preferences may allow to integrate the training pipeline in robot programming software. Finally, exploring more complex manipulation scenarios and multi-agent learning setups could extend the applicability of our method to a broader range of real-world tasks. Our findings highlight the potential of LLMs in accelerating RL-based robotic learning, paving the way for more autonomous and efficient robot training paradigms.
Acknowledgments
These experiments were carried out in the WASP Research Arena (WARA)-Robotics, hosted by ABB Corporate Research Center in Västerås, Sweden and financially supported by the Wallenberg AI, Autonomous Systems, and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix
VI-A GPT-4 prompts for reward generation
@@
The task required for the GPT-4 model is a code generation task, in this appendix, we report the prompt structure we used. You are a reward engineer trying to write reward functions to solve reinforcement learning tasks as effective as possible. Your goal is to write a dense reward function for the environment that will help the agent learn the task described in text. Your reward function should use useful variables from the environment. Write a python class that satisfies the following requirements: -The new class must inherit from the class defined in the code snipped included after the requirements. -The new class must not redefine the constructor -The new class must only redefine the reward_fun and _get_info methods to encode the following task for the environment: TASK DESCRIPTION -The reward_fun method must define a dictionary named reward_dict, reward_dict must contain all the values of the reward components, excluding the reward for solving the task. Then the method must conclude the following code: total_bonuses = sum([rewards_dict[k] if rewards_dict[k] >>> 0 else 0 for k in rewards_dict.keys()]) total_bonuses = total_bonuses if total_bonuses > 0 else 1 task_solved_reward = 10 * self._max_episode_steps * total_bonuses * info[’task_solved’] total_shaping = sum([rewards_dict[k] for k in rewards_dict.keys()]) reward = total_shaping + task_solved_reward rewards_dict[’task_solved_reward’] = task_solved_reward return reward, rewards_dict The _get_info method must return a dict that contains a boolean field task_solved, that is computed according to the task description provided. If the prompt requires, the new class can implement a termination_condition method, that returns True if a failure condition is detected. Not all tasks require this failure condition to be implemented. BASE CLASS CODE Here is some advice on how to design reward functions: -The reward function should implement a shaping that clearly guides the policy towards the goal -You can give a rewarding bonus if the task is partially solved. -If the environment has a termination condition that can halt an episode before reaching the goal area or configuration, consider using positive shaping terms in the reward function. -The reward components should not be too disproportionate. -To incentive to get close to an object you should reward the decrease of distance and the contact with said objects, if you want to avoid touching another object just give negative rewards if that is touched. -If you want to grasp an object and/or not to drop it you should reward contacts with the gripper fingers and the object -To reward lifting of objects, you can assume that the 3rd dimension of position vectors is the vertical axis, which is oriented towards the ceiling. You can use +height as reward shaping for lifting tasks. -Penalties for unwanted actions should be very small in absolute value, compared to positive rewards. Here is one example of reward function for a cube picking task: cube_pose = np.array(sim_utils.getObjPose(self.cube_id)) end_effector_left_pose = np.array(sim_utils.getLinkPose(self.robot_id, self.end_effector_left_idx)) distance = np.linalg.norm(np.array(cube_pose[:3]) - np.array(end_effector_left_pose[:3])) finger_1, finger_2, total_contact_force = self.check_fingers_touching(self.cube_id) rewards_dict = { "distance_to_cube_reward": -distance, # encourage getting close to the target cube "finger_contacts_reward": 1 * finger_1 + 1 * finger_2, # encourage interaction with the target cube "cube_height_reward": 1000 * (cube_pose[2] - 0.02) }
@@
VI-B Robotic manipulation task descriptions and generated rewards
VI-B1 CubeSlide
@@
17 total_bonuses = total_bonuses if total_bonuses > 0 else 1
@@ VI-B2 CubePick
30 total_bonuses = total_bonuses if total_bonuses > 0 else 1
@@ VI-B3 CubeStack
18 "distance_to_red_cube_reward": -10*xy_distance, # manual
19 "height_diff_reward": -10*height_diff, # manual
@@
25 total_bonuses = total_bonuses if total_bonuses > 0 else 1
@@ VI-B4 CubesTowerSlide
18 total_bonuses = total_bonuses if total_bonuses > 0 else 1
@@ VI-B5 VialGrasp
25 total_bonuses = total_bonuses if total_bonuses > 0 else 1
@@ VI-B6 CentrifugeInsertion
16 total_bonuses = total_bonuses if total_bonuses > 0 else 1
@@ VI-B7 BimanualPenInTape
26 total_bonuses = total_bonuses if total_bonuses > 0 else 1
@@ VI-B8 CubeInCup
24 total_bonuses = total_bonuses if total_bonuses > 0 else 1
@@ VI-B9 BimanualBoxLift
24 total_bonuses = total_bonuses if total_bonuses > 0 else 1
@@
21 total_bonuses = total_bonuses if total_bonuses > 0 else 1

### v4

@@ VI-A Pseudocode
VI-B GPT-4 prompts for reward generation
@@
The task required for the GPT-4 model is a code generation task, in this appendix, we report the prompt structure we used. You are a reward engineer trying to write reward functions to solve reinforcement learning tasks as effective as possible. Your goal is to write a dense reward function for the environment that will help the agent learn the task described in text. Your reward function should use useful variables from the environment. Write a python class that satisfies the following requirements: -The new class must inherit from the class defined in the code snipped included after the requirements. -The new class must not redefine the constructor -The new class must only redefine the reward_fun and _get_info methods to encode the following task for the environment: TASK DESCRIPTION -The reward_fun method must define a dictionary named reward_dict, reward_dict must contain all the values of the reward components, excluding the reward for solving the task. Then the method must conclude the following code: total_bonuses = sum([rewards_dict[k] if rewards_dict[k] >>> 0 else 0 for k in rewards_dict.keys()]) total_bonuses = max(total_bonuses, 1) task_solved_reward = 10 * self._max_episode_steps * total_bonuses * info[’task_solved’] total_shaping = sum([rewards_dict[k] for k in rewards_dict.keys()]) reward = total_shaping + task_solved_reward rewards_dict[’task_solved_reward’] = task_solved_reward return reward, rewards_dict The _get_info method must return a dict that contains a boolean field task_solved, that is computed according to the task description provided. If the prompt requires, the new class can implement a termination_condition method, that returns True if a failure condition is detected. Not all tasks require this failure condition to be implemented. BASE CLASS CODE Here is some advice on how to design reward functions: -The reward function should implement a shaping that clearly guides the policy towards the goal -You can give a rewarding bonus if the task is partially solved. -If the environment has a termination condition that can halt an episode before reaching the goal area or configuration, consider using positive shaping terms in the reward function. -The reward components should not be too disproportionate. -To incentive to get close to an object you should reward the decrease of distance and the contact with said objects, if you want to avoid touching another object just give negative rewards if that is touched. -If you want to grasp an object and/or not to drop it you should reward contacts with the gripper fingers and the object -To reward lifting of objects, you can assume that the 3rd dimension of position vectors is the vertical axis, which is oriented towards the ceiling. You can use +height as reward shaping for lifting tasks. -Penalties for unwanted actions should be very small in absolute value, compared to positive rewards. Here is one example of reward function for a cube picking task: cube_pose = np.array(sim_utils.getObjPose(self.cube_id)) end_effector_left_pose = np.array(sim_utils.getLinkPose(self.robot_id, self.end_effector_left_idx)) distance = np.linalg.norm(np.array(cube_pose[:3]) - np.array(end_effector_left_pose[:3])) finger_1, finger_2, total_contact_force = self.check_fingers_touching(self.cube_id) rewards_dict = { "distance_to_cube_reward": -distance, # encourage getting close to the target cube "finger_contacts_reward": 1 * finger_1 + 1 * finger_2, # encourage interaction with the target cube "cube_height_reward": 1000 * (cube_pose[2] - 0.02) }
@@
VI-C Benchmark Environments Task descriptions
VI-C1 Grasp and Lift
Task: The agent mush reach the cube position. Then grasp and lift the cube. Consider the task solved when the cube is at 0.5 height. There is no failure condition.
VI-C2 Grasp and Slide
Task: The agent mush reach the cube position. Then grasp the cube. Finally, move the cube to the right (positive x direction). Consider the task solved when the cube x coordinate is $>$ 0.99. There is no failure condition.
VI-C3 Place
Task: The agent is holding the cube. Move the cube to (0, 0). Consider the task solved when the cube is at 0.05 distance from (0, 0) and is grasped. Consider the task failed when the agent doesn’t grasp the cube.
VI-C4 Push Cube
Task: The agent and the cube are on a narrow table. The agent mush reach the cube position. Then the agent must push the cube towards the positive x direction, without dropping the cube from the table. Consider the task solved when the cube is at x position $>$ 0.5. Consider the task failed when the cube is at z height $<$ table_height.
VI-D Example of generated reward ”fixed” by the proposed formalization
As we argue in the paper, the weight imbalance issue is a major issue for reward generation with LLMs such as GPT-4. In the included code, we have multiple examples of this being “fixed“ by the terminal reward defined in this formulation. In the following, we consider one example of a reward function generated by ARCHIE with GPT-4, for the cube push task (Section V-A):
$r(s_{t},a_{t})=-d+10\cdot contact+x_{t},$ (9)
where $d$ is the distance between the two cubes, $contact$ is a boolean variable which is true when the cubes are in contact (part of the observation), and $x_{t}$ is the x coordinate of the blue cube. This is a typical example of reward with correct structure and unbalanced weights, since $d$ and $x_{t}$ are both in the order of $10^{0}$, while the weight of $contact$ is one order of magnitude greater. It is similar to the case in Figure 6 of the manuscript. With the terminal term introduced in the novel formalization, the reward then becomes:
$r(s_{t},a_{t},s_{t+1})=-d+10\cdot contact+x_{t}+\underbrace{R_{F}(s_{t},a_{t})\Phi(s_{t+1})}_{\text{terminal term}},$ (10)
with
$\Phi(s_{t})=\begin{cases}1\quad\text{if }x_{t}>table\_lenght\\ 0\quad\text{otherwise}.\end{cases}$ (11)
and $R_{F}(s_{t},a_{t})$ in the order of $10^{2}$ when $\Phi(s_{t+1})=1$.
This example of reward was generated by GPT-4 in this Python code:
def _get_info(self):
    cube_pos, _ = p.getBasePositionAndOrientation(self.cube2_id)
    task_solved = cube_pos[0] > self.table_lenght and cube_pos[2] >= self.table_height
    return {’task_solved’: task_solved}
def reward_fun(self, observation, action):
    agent_pos, _ = p.getBasePositionAndOrientation(self.agent)
    cube_pos, _ = p.getBasePositionAndOrientation(self.cube2_id)
    rewards_dict = {
    # encourage proximity
        "distance_to_cube_reward": -np.linalg.norm(np.array(cube_pos[:3]) - np.array(agent_pos[:3])),
        "contact_reward": 10 * self.check_contact(), # reward contact with the cube
        "x_direction_push_reward": cube_pos[0], # reward pushing in the x-direction
    total_bonuses = sum([val if val > 0 else 0 for val in rewards_dict.values()])
    total_bonuses = max(total_bonuses, 1)
    info = self._get_info()
    total_shaping = sum(rewards_dict.values())
    task_solved_reward = 10 * self._max_episode_steps * total_bonuses * info[’task_solved’]
    reward = total_shaping + task_solved_reward
    rewards_dict[’task_solved_reward’] = task_solved_reward
    return reward, rewards_dict
In Section V-A, Figure 6 shows the learning curves of SAC agents trained with this reward and with its version without the terminal term. Note Agents trained with the terminal reward term have episode termination enabled when the task is solved. Instead, the agents trained without the terminal term, i.e. the reward completely generated by GPT-4, do not have episode termination on completion enabled
VI-E Reward weights unbalance example
Referring to the reward generation with LLMs described in Section IV-B, with the available context, the shaping terms generated by the LLM are generally coherent with the task requirements. Nevertheless, there is no guarantee that the different generated terms are numerically coherent with each other or balanced, as discussed in other works. As an example, for a pushing task, where the agent is required to move an object from a starting position to a target position, one possible reward function that LLMs are capable of generating is:
where $d$ is the distance between the object and the target position, while $b\in\mathbb{N^{+}}$ is a bonus for touching the object.
The bonus term $b$ can help guide exploration, but even simple rewards require tuning. If $b$ is too large relative to $d$, the reward landscape becomes too flat, leading to suboptimal learning. We illustrate this with a 2D example where a point-agent moves toward the origin. Using the reward from eq. 12, we compare SAC training with $b=10$ and $b=1$. As shown in Figure 6, a large $b$ flattens the reward landscape, hindering learning, while a smaller $b$ offers clearer guidance and better performance. This highlights that even well-designed rewards can fail without proper weighting.
Figure 6: Top: Rewards landscapes for two example rewards. Bottom: statistics of success rates of the SAC agents trained with the two rewards.
VI-F Robotic manipulation task descriptions and generated rewards
VI-F1 CubeSlide
@@
17 total_bonuses = max(total_bonuses, 1)
@@ VI-F2 CubePick
30 total_bonuses = max(total_bonuses, 1)
@@ VI-F3 CubeStack
18 "distance_to_red_cube_reward": -10*xy_distance,
19 "height_diff_reward": -10*height_diff,
@@
25 total_bonuses = max(total_bonuses, 1)
@@ VI-F4 CubesTowerSlide
18 total_bonuses = max(total_bonuses, 1)
@@ VI-F5 VialGrasp
25 total_bonuses = max(total_bonuses, 1)
@@ VI-F6 CentrifugeInsertion
16 total_bonuses = max(total_bonuses, 1)
@@ VI-F7 BimanualPenInTape
26 total_bonuses = max(total_bonuses, 1)
@@ VI-F8 CubeInCup
24 total_bonuses = max(total_bonuses, 1)
@@ VI-F9 BimanualBoxLift
24 total_bonuses = max(total_bonuses, 1)
@@
21 total_bonuses = max(total_bonuses, 1)